OpenStudy (anonymous):

Help? The sum of the digits of a certain two digit number is 10. When you reverse the digits you increased the number by 54. what is the number? Anyone help me with these steps?

OpenStudy (anonymous):

x + y = 10 10x + y = (x + 10y) + 54

OpenStudy (anonymous):

i'm confused.. can you explain the second equation?

OpenStudy (anonymous):

If you consider x and y to be digits, then 10x will put the x in the "10"s. Same for the other side of the equation. It's just reversing. That equation comes down to: 9x - 9y = 54 x - y = 6

OpenStudy (anonymous):

Now, you only have to solve x + y = 10 and x - y = 6 Two simultaneous equations in 2 variables.

OpenStudy (mathstudent55):

let u = units digit in original number let t = tens digit in original number Original number: t + u = 10 New number 10u + t = 10t + u + 54 Solve the equations as a system of equations.

OpenStudy (anonymous):

You can take the 2nd equation and call it x = 6 - y and substitute it into the first equation and get your value for "y".

OpenStudy (anonymous):

So, (6 + y) + y = 10 6 + 2y = 10 -> y = 2 So, x = 8 Take a look at numbers 28 and 82 and see if they come up with the difference you are looking for.

OpenStudy (anonymous):

82 - 28 = 54 So, these are the 2 numbers that when reversed have a diference of 54 and the digits themselves add up to 10.

OpenStudy (anonymous):

All good now @Artchicky ?

OpenStudy (anonymous):

not really. Still trying to understand the first part.

OpenStudy (anonymous):

When you think of a digit, it will have a "place" in the number. If you take one of the digits, say "x", and multiply it 10, so that it is 10x, it will look like x_ where the "-" is the "units" place and the "x" occupies the "10"s position. It's like x-ty Example, if your digit is 8 and you multiply it by 10, you get 80 or eighty.

OpenStudy (anonymous):

Ohh ok i get it

OpenStudy (anonymous):

Cool!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @Artchicky

OpenStudy (anonymous):

Thank you for your help and no problem!

OpenStudy (anonymous):

You're very welcome! Nice working with you!